Suppose 90-day investments in Britain have a 6% annualized return and a 1.5% quarterly (90-day) return. In the U.S., 90-day investments of similar risk have a 4% annualized return and a 1% quarterly (90-day) return. In the 90-day forward market, 1 British pound equals $1.65. If interest rate parity holds, what is the spot exchange rate ($/£)?





1
Expert's answer
2021-01-11T07:45:00-0500

1 pound = $1.6582, $1.65 x (1 + 0.015) / (1 + 0.01) = 1.67475 / 1.01 = $1.6582